### v1

# Multi-Agent Reinforcement Learning with Control-Theoretic Safety Guarantees for Dynamic Network Bridging

Abstract
Addressing complex cooperative tasks in safety-critical environments poses significant challenges for Multi-Agent Systems, especially under conditions of partial observability. This work introduces a hybrid approach that integrates Multi-Agent Reinforcement Learning with control-theoretic methods to ensure safe and efficient distributed strategies. Our contributions include a novel setpoint update algorithm that dynamically adjusts agents’ positions to preserve safety conditions without compromising the mission’s objectives. Through experimental validation, we demonstrate significant advantages over conventional MARL strategies, achieving comparable task performance with zero safety violations. Our findings indicate that integrating safe control with learning approaches not only enhances safety compliance but also achieves good performance in mission objectives.
I Introduction
Despite the promising performance of Multi-Agent Reinforcement Learning (MARL) in simulations, its application in real-world, safety-critical scenarios raises significant concerns. A key challenge is collision avoidance, especially in applications like autonomous driving, where failure to prevent collisions could result in severe consequences. Moreover, the inherent nature of Multi-agent Systems (MAS) often involves partial observability, further increasing the complexity of these problems by limiting the information available to each agent about the environment and the states of other agents.
This challenge highlights a fundamental limitation of all Reinfocement Learning (RL)-based approaches: while agents learn to optimize predefined reward functions, their behavior lacks guarantees and remains inherently unpredictable. The sole reliance on MARL’s reward functions to ensure safety is insufficient [1]. Even with extensive training and monitoring of an agent’s performance based on expected rewards, or other performance measures, it is impossible to exhaustively examine all potential scenarios in which the agent might fail to act safely or predictably.
In ensuring safety among agents, various control-theoretic approaches and hybrid learning approaches have been developed, ranging from Model Predictive Control (MPC) [2] to Reference Governor (RG) techniques [3], and from Control Barrier Functions (CBFs) [4] to Reachability Analysis (RA) methods [5]. In this work, we focus on developing a hybrid learning approach built upon the method used in [6], that can learn to achieve complex objectives while ensuring safety guarantees in distributed MAS.
To demonstrate and evaluate the effectiveness of our hybrid approach we consider the cooperative task of Dynamic Network Bridging [7] (Fig. 1). In this task the goal is to establish and maintain a connection between two moving targets A and B in a 2D plane, relying on a swarm of $N$ agents with limited observation of the environment. The agents must form an ad-hoc mobile network that establishes a communication path between the two targets as they move, dynamically adjusting their positions to ensure an uninterrupted multi-hop communication link.
Figure 1: Decentralized swarm coordination for dynamic network bridging
In this paper we introduce a hybrid approach combining MARL and control-theoretic methods, accomplishing task objectives while providing safety guarantees. Our key contributions are: 1) A decentralized control framework that is MARL compatible and restricts the effect of each agent’s movement updates to only its one-hop neighbors, enabling efficient local coordination; 2) An algorithm that updates setpoints while preserving safety conditions through communication with affected neighbors; 3) An analytical computationally tractable condition verifying potential safety violations during updates. Overall our approach provides effective coordination among agents while ensuring safety constraints.
II Preliminaries
In this section, we present the foundational theoretical elements, drawing from both control theory and MARL, that form the basis of our proposed hybrid approach.
Consider a linear time invariant (LTI) system expressed in the state space form
$\dot{x}=Ax+Bu$ (1)
where $x\in\mathbb{R}^{n}$ is the state vector with $n$ components, and $u\in\mathbb{R}^{p}$ is the input vector with $p$ components. Thus, the dynamical matrices are $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times p}$. If the pair $(A,B)$ is controllable, there exists a state feedback controller:
$u=-Kx$ (2)
such that, the origin ($x=0$) of the resulting closed-loop (C-L) system:
$\dot{x}=(A-BK)x=A_{f}x$ (3)
is an asymptotically stable equilibrium point (EP): $x(t)\rightarrow 0\quad\text{as}\quad t\rightarrow+\infty$. Hence, the matrix $A_{f}$ is Hurwitz, i.e. all the eigenvalues have a negative real part [8]. The problem to design (2) that makes (1) asymptotically stable is called stabilization problem.
II-A Tracking Control Problem
In the case the problem requires to move the system (1) following a specific trajectory, the problem is called tracking control. This problem can be solved by generating a sequence of EPs [9] that we call setpoints, denoted by $x_{sp}$. The C-L system (3) can be represented as:
$\dot{x}=A_{f}(x-x_{sp})$ (4)
where the control input (2) becomes $u=-K(x-x_{sp}).$
II-B Lyapunov’s Theory
Since $A_{f}$ is Hurwitz, there exists a symmetric positive definite matrix $P>0$ such that:
$A_{f}^{T}P+PA_{f}=-Q$ (5)
for a given symmetric positive definite matrix $Q>0$. Equation (5) is called Lyapunov’s equation. Defining the $P$-norm as:
$\lVert x-x_{sp}\rVert_{P}\triangleq\sqrt{(x-x_{sp})^{T}P(x-x_{sp})}$ (6)
the $P$-norm squared is a Lyapunov function $V(x)=\lVert x-x_{sp}\rVert_{P}^{2}$, which is always positive except at $x_{sp}$ where it is zero. For all trajectories $x(t)$ starting within the ellipsoid:
$\mathcal{E}_{c}(x_{sp})=\{x\in\mathbb{R}^{n}\,|\,\lVert x-x_{sp}\rVert_{P}^{2}\leq c\},$ (7)
where $c>0$, $V(x(t))$ is monotonically decreasing. Therefore, the trajectory $x(t)$ never leaves $\mathcal{E}_{c}$, which is called a positively invariant set [8].
Remark 1
In the case of a nonlinear dynamical system described by $\dot{x}=f(x,u)$ where $u=0$, $x=0$ is an equilibrium point for the system $f(0,0)=0$. The control input (2) can be designed by considering the linearized system (1) around the equilibrium point. Therefore, the closed-loop (CL) system (3) can be used to solve Lyapunov’s equation (5), which is used to compute the positively invariant set $\mathcal{E}_{c}(x_{sp})$ (7). While for LTI systems, $\mathcal{E}_{c}(x_{sp})$ is positively invariant for any positive $c\in\mathbb{R}$, for nonlinear systems, we need to verify that condition in a neighborhood of the equilibrium point $x_{sp}$ [8]. Furthermore, for nonlinear systems, $x_{sp}$ has to satisfy $f(x_{sp},0)=0$.
II-C 6-DOF UAV
In this work, we consider a 6-DOF UAV as the agent, described by a state vector of 12 components:
$x=\begin{bmatrix}\dot{p}_{x},&p_{x},&\dot{p}_{y},&p_{y},&\dot{p}_{z},&p_{z},&\dot{\phi},&\phi,&\dot{\theta},\\ \theta,&\dot{\psi},&\psi\end{bmatrix}^{T}$ (8)
where $p_{x}$, $p_{y}$, $p_{z}$, and their derivatives represent linear position and velocity, respectively. Similarly, $\phi$, $\theta$, $\psi$, and their derivatives represent the Euler angles (roll, pitch, and yaw) and their angular velocities, respectively. This model is consistent with the one used in [9] and references therein. The controller used is given by (2), obtained by linearizing the quadrotor model around an equilibrium point. For this case, the equilibrium point is indicated with $x_{sp}$ which has the same structure of (8) with all components equal to zero except for $p_{x}$, $p_{y}$, $p_{z}$.
II-D Multi-Agent Reinforcement Learning
In a typical multi-agent cooperative scenario, $N$ agents interact within an environment to complete a task. Noisy and limited sensors may prevent each agent from observing the full state of the environment, allowing access only to partial observations. With partial observability, an agent no longer knows the true state of the environment, but instead needs to maintain a belief state - a probability distribution over states estimated from its observation history - which it uses to select actions. Decision-making processes in such contexts can be formulated as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) [10, 11].
A Dec-POMDP is defined by a tuple $(S,\mathbf{A},\mathcal{O},T,R,\Omega,\gamma,N)$ where $S$ is a finite set of states of the environment, $\mathbf{A}$ is a set of joint action spaces, and $N$ is the number of agents, $T$ is the state transition probability function which gives the probability of transitioning from state $s$ to state $s^{\prime}$ when joint action $\mathbf{a}=(a_{1},\cdots,a_{N})$ is taken; $R$ is the joint reward function that maps states and joint actions to real numbers and is used to specify the goal of the agents; $\Omega$ is a set of joint observations $O_{i}$, where $O_{i}$ is the observation set for agent $i$; $\mathcal{O}$ is the joint observation probability function, which gives the probability of receiving joint observation $\mathbf{o}=(o_{1},\cdots,o_{N})$ after taking joint action $\mathbf{a}$ and ending up in state $s^{\prime}$; and $\gamma$ is the reward discount factor, $0\leq\gamma\leq 1$.
In a Dec-POMDP, agents need to collaborate and coordinate their actions based on their individual observations to maximize the joint reward function. Solving a Dec-POMDP optimally is computationally intractable, as the policy space grows exponentially with the number of agents and the planning horizon. Therefore, various approximate solution methods [12, 13] have been used, such as heuristic search, dynamic programming, value or policy-based approximators with neural networks etc. The problem facing the team is to find the optimal joint policy, i.e. a combination of individual agent policies that produces behavior that maximizes the team’s expected reward.
III Problem Formulation
We address the challenge of dynamically forming and sustaining a connection between two moving targets, A and B, in a 2D space using a swarm of $N$ agents [7]. These agents must collaboratively adjust their positions to maintain a continuous multi-hop link between the targets, operating under decentralized control and limited local perception, while ensuring they avoid collisions and respect physical boundaries. We consider all the agents as nonlinear systems that can be reduced into the form (1) controlled by (2) around the equilibrium point $x_{sp,i}$ [9]:
$\dot{x}_{i}=A_{f}(x_{i}-x_{sp,i})=(A-BK)(x_{i}-x_{sp,i})$ (9)
for $i=1,\ldots,N$. In this case, we consider $N$ agents described with the same model, but in general, our approach can be generalized to the case of different dynamics and multidimensional scenarios.
III-A One-hop Communication
We assume that for each agent, there is a ball of radius $r$ in the 2- or 3-dimensional Euclidean space centered at the agent’s current linear position $p_{i}(t)=[p_{x,i}(t),p_{y,i}(t),p_{z,i}(t)]^{T}$:
$\mathcal{B}_{r}(p_{i}(t))=\left\{p\in\mathbb{R}^{3}:\|p-p_{i}(t)\|^{2}\leq r^{2}\right\}$
Given two agents $i$ and $j$ with $i\neq j$, they can communicate if and only if
$\mathcal{B}_{r}(p_{i}(t))\cap\mathcal{B}_{r}(p_{j}(t))\neq\emptyset$ (10)
In our setup, since $p_{z,i}$ is fixed, the one-hop communication region is described by a 2-D ball. Note that, we use a different notation when we refer to the 2-D communication space to avoid any confusion between the quadrotor $x$-axis and the state space $x$.
III-B Safety
Definition 1
Given two agents $i$ and $j$, where $x_{sp,i}\neq x_{sp,j}$, $x_{i}(t_{0})\in\mathcal{E}_{c}(x_{sp,i})$, and $x_{j}(t_{0})\in\mathcal{E}_{c}(x_{sp,j})$. Agents $i$ and $j$ are safe with respect to each other if
$\mathcal{E}_{c}(x_{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})=\emptyset.$ (11)
The above definition is derived by the positively invariant property of (7), since the initial conditions of each agent are inside their respective ellipsoids, then $x_{i}(t)\in\mathcal{E}_{c}(x_{sp,i})$ and $x_{j}(t)\in\mathcal{E}_{c}(x_{sp,j})$ for $t\rightarrow+\infty$. This means that the trajectory of each agent cannot intersect generating the risk of possible collisions.
In order to guarantee safety in our setup, we need the following assumption:
Assumption 1
The projection of $\mathcal{E}_{c}(x_{sp,i})$ into the 2-D space formed by the components $p_{x}$ and $p_{y}$ must be contained within the 2-dimensional ball $\mathcal{B}_{r}(p_{i}(t))$, where $p_{i}(t)=[p_{x,i}(t),p_{y,i}(t)]$ represents the actual position coordinates ($p_{x,i}(t),p_{y,i}(t)$) of agent $i$ for any state $x_{i}\in\mathcal{E}_{c}(x_{sp,i})$.
III-C Cooperative MARL for Dynamic Network Bridging
Each agent $i$ has a local observation $o_{i}$ comprising the structure of its neighborhood and features representing its neighbors at certain time-step $t$. Such features include the node ID, current coordinates, action taken, and coordinates of the targets A and B to be connected. At the beginning of every time-step, agents make decentralized decisions, generating target points $w$, utilizing their local observation to choose the direction of their next movement. This is encoded in a two-dimensional action space, with each dimension having three options: move forward, move backward, or hold along the corresponding $p_{x}$ or $p_{y}$ axis. Given the direction chosen by the agents, a new target point is calculated for each agent using a fixed offset equal for both the $p_{x}$ and $p_{y}$ axes. Finally, agents are rewarded using a reward function that motivates agents to form a connected network bridging the targets. Such reward signal combines three components:
Base Connectivity Reward: The base reward, $R_{\text{base}}$, encourages larger connected components:
$R_{\text{base}}(s)=\frac{|C_{\text{max}}(s)|}{|\mathcal{V}|},$ (12)
where $|C_{\text{max}}(s)|$ represents the size of the largest connected component in state $s$, and $|\mathcal{V}|$ is the total number of entities.
Centroid Distance Penalty: The centroid distance penalty, $P_{\text{cent}}$, is based on agents’ positions relative to the targets, penalizing them based on the Euclidean distance between the centroid of the agents’ positions and the central point between the targets.
Target Path Bonus: A bonus, $B_{\text{path}}=100$, is awarded if a path exists between the two targets.
The overall reward combines these three elements:
$R(s,a)=\begin{cases}B_{\text{path}}(s)&\mbox{if }\exists\mathrm{path}(T_{1},T_{2});\par\\ R_{\text{base}}(s)-P_{\text{cent}}(s),&\mbox{otherwise}.\end{cases}$ (13)
We also include a physics simulator, modeling the continuous non-linear dynamics of the agents (2), and a safe tracking control system proposed in the next section (Algorithm 1).
IV Safe Tracking Control
The safety tracking control algorithm that we propose here consists of two steps, the first one is the setpoint update of the agent $i$ that guarantees that its current position belongs to $\mathcal{E}_{c}$ of the new setpoint. This condition is fundamental for the second step where the safety condition (11) is checked for all agents within the one-hop communication ball.
Let us start with the first step which has been borrowed from [6]. For each agent, we select two scalars $c$ and $s$ such that $0<s<c$ therefore $\mathcal{E}_{s}(x_{sp,i})\subset\mathcal{E}_{c}(x_{sp,i})$.
Definition 2
We can say that the agent has reached the setpoint $x_{sp,i}$ if and only if $x_{i}(t)\in\mathcal{E}_{s}(x_{sp,i})$.
Assuming that the agent $i$ is moving from the target point $w_{i}$ to the target point $w_{i+1}$, generated by the MARL, the idea is to generate a sequence of setpoints on the line that connects the two target points. Once the state of agent $i$, $x_{i}$ reached the setpoint $x_{sp,i}$, a new setpoint $x^{\prime}_{sp,i}$ is generated following the rule
$x^{\prime}_{sp,i}=x_{sp,i}+(\sqrt{c}-\sqrt{s})v$ (14)
where $v\in\mathbb{R}^{n}$ with $\|v\|_{2}=1$, indicating the direction of the line connecting the two target points. As showed in [6]
$x_{i}(t)\in\mathcal{E}_{s}(x_{sp,i})\Rightarrow x_{i}(t)\in\mathcal{E}_{c}(x^{\prime}_{sp,i}).$ (15)
Fig. 2 illustrates the process of updating the setpoint, presented in 2-D space for clearer visualization, though it actually corresponds to a state space with n=12 components, as applicable to quadrotors.
Figure 2: Setpoint update for agent $i$ in the state space. Note that the target points $w_{i}$ and $w_{i+1}$ have the same structure of the setpoints $x_{sp,i}$.
Thanks to (14) and Assumption 1, the projection of $\mathcal{E}_{c}(x^{\prime}_{sp,i})$ is contained in $\mathcal{B}(p_{i}(t))$.
Proposition 1
Considering agent $i$ moving from target point $w_{i}$ to $w_{i+1}$, and updating the setpoint $x_{sp,i}$ every time $x_{i}(t)\in\mathcal{E}_{s}(x_{sp,i})$ by using (14). Considering also Assumption 1 true, then the safety condition (11) can be violated only with all the agents $j\neq i$ that satisfy (10) (i.e. within the one-hop communication range).
Proof:
From Assumption 1, the ball $\mathcal{B}_{r}(p_{i}(t))$ encompasses the projection onto the 2-D space of $\mathcal{E}_{c}(x{sp,i})$ for any current state $x_{i}(t)\in\mathcal{E}_{c}(x_{sp,i})$. Utilizing (14) to adjust the setpoint ensures that $x_{i}(t)$ remains within $\mathcal{E}_{c}(x^{\prime}_{sp,i})$, signifying that its 2-D space projection still falls within $\mathcal{B}_{r}(p_{i}(t))$. Thus, should the updated setpoint $x^{\prime}_{sp,i}$ contravene (11), such a violation would only occur concerning agents situated within the one-hop communication range. ∎
Proposition 1 highlights a critical aspect of setpoint updates: setpoint updates only impact safety for agents within one-hop range. This ensures agent $i$ can communicate with affected neighbors before executing updates. This allows coordinating to preemptively address any potential safety concerns from the new setpoint through direct communication with impacted agents.
Building upon Proposition 1, let’s define $\mathcal{D}_{i}(t)$ as the set comprising all agents $j\neq i$ that are within the one-hop communication range of agent $i$ at time $t$. The safe setpoint update algorithm for each agent $i$ is given in Algorithm 1. Once agent $i$’s state $x_{i}(t)$ reaches its current setpoint $x_{sp,i}$ (as per Definition 2), it computes a new target setpoint $x^{\prime}_{sp,i}$. Before updating, agent $i$ checks for any possible safety violations by evaluating the safety condition (11) for all agents in $\mathcal{D}_{i}(t)$. If no violation is detected, $x_{sp,i}$ is updated to the new $x^{\prime}_{sp,i}$; otherwise, the previous setpoint is retained.
Input: Agent $i$ with setpoint $x_{sp,i}$, $\mathcal{D}_{i}(t)$
Output: Updated setpoint $x_{sp,i}$
safety_violation $\leftarrow$ 0;
if $x_{i}(t)\in\mathcal{E}_{s}(x_{sp,i})$ then
       Compute $x^{\prime}_{sp}\leftarrow$ (14);
       for $j\in\mathcal{D}_{i}(t)$ do
             if $\mathcal{E}_{c}(x_{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})\neq\emptyset$  then
                   safety_violation $\leftarrow 1$;
                   break;
             end if
       end for
       if safety_violation ==absent$=$ = 00 then
             $x_{sp,i}\leftarrow x^{\prime}_{sp,i}$;
       end if
end if
Algorithm 1 Setpoint Update Algorithm
By applying Algorithm 1 to each agent, we ensure that the safety condition (11) is satisfied throughout the system. This is guaranteed by the fact that an agent can only update its setpoint after verifying that the new target does not violate safety for any agents within its one-hop communication range, as stated in Proposition 1. The proof of safety relies on the following assumptions:
Assumption 2
The initial setpoints of all agents satisfy the safety condition (11).
Assumption 3
During the setpoint update for agent $i$, all agents $j\in\mathcal{D}_{i}(t)$ cannot change their setpoints $x_{sp,j}$.
Proposition 2
Considering the multi-agent system described by (9) and Assumptions 1, 2 and 3, then applying Algorithm 1 to each agent guarantees that the safety condition (11) is satisfied for all agents.
Proof:
By Assumption 2, the initial setpoints of all agents satisfy the safety condition (11). Whenever an agent $i$ updates its setpoint using Algorithm 1, it checks if the new target setpoint $x_{sp,i}^{\prime}$ violates the safety condition for any agent $j\in\mathcal{D}_{i}(t)$. If a violation is detected, the setpoint is not updated, ensuring that the current safe solution is maintained. Proposition 1 guarantees that safety is only impacted for agents within the one-hop communication range, which agent $i$ can directly communicate with before executing the update. Furthermore, Assumption 3 prevents agents in $\mathcal{D}_{i}(t)$ from changing their setpoints during agent $i$ ’s update, ensuring a consistent safety evaluation. Therefore, by applying Algorithm 1 to each agent, the safety condition (11) is upheld throughout the system’s operation. ∎
We now need to verify when the safety condition (11) is violated. Our objective is to demonstrate that if (11) is violated, then there exists a point on the line connecting $x_{sp,i}$ and $x_{sp,j}$, inside the overlap region of two ellipsoids $\mathcal{E}_{c}(x_{sp,i})$ and $\mathcal{E}_{c}(x_{sp,j})$ (Fig. 3). This overlap region, denoted by $\mathcal{E}_{c}(x{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})$, satisfies the following condition:
$\lambda(x-x_{sp,i})^{T}P(x-x_{sp,i})+(\lambda-1)(x-x_{sp,j})^{T}P(x-x_{sp,j})\leq c$ (16)
for $0\leq\lambda\leq 1$ with $x_{sp,i}\neq x_{sp,j}$.
Proposition 3
Let us consider the ellipsoids $\mathcal{E}_{c}(x_{sp,i})$ and $\mathcal{E}_{c}(x_{sp,j})$ defined as in (7). The set of points $x$ satisfying (16) for all $\lambda\in[0,1]$ is either empty, a single point, or an ellipsoid:
^$\hat{\mathcal{E}}_{K_{\lambda}}(m_{\lambda})=\left\{x\in\mathbb{R}^{n}:(x-m_{\lambda})^{T}P^{-1}(x-m_{\lambda})\leq K_{\lambda}\right\}$ (17)
where
$m_{\lambda}$ $=\lambda x_{sp,i}+(1-\lambda)x_{sp,j}$
$K_{\lambda}$ $=1-\lambda(1-\lambda)(x_{sp,j}-x_{sp,i})^{T}P(x_{sp,j}-x_{sp,i}).$
Proof:
See [14]. ∎
Proposition 3 shows that if there is an ellipsoid violation, then the point $m_{\lambda}$ belongs to the intersection Fig. 3:
$\mathcal{E}_{c}(x_{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})\neq\emptyset\rightarrow m_{\lambda}\in\mathcal{E}_{c}(x_{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})$
Figure 3: Explaination of Proposition 3: $m_{\lambda}$ is on the segment that connects the two setpoints.
Therefore, to test if the two ellipsoids intersect we need to find if there exists at least one point on the segment joining $x_{sp,i}$ and $x_{sp,j}$ that satisfies (16). To do so we consider both ellipsoids
$\left\{\ \begin{array}[]{l}(x-x_{sp,i})^{T}P(x-x_{sp,i})\leq\rho\\ (x-x_{sp,j})^{T}P(x-x_{sp,j})\leq\rho\\ \end{array}\right.$ (18)
We replace $x$ with $m_{\lambda}$ since we are checking only a point in the segment joining the two setpoints. Defining $d\triangleq x_{sp,j}-x_{sp,i}$, (18) can be rewritten as
$\left\{\ \begin{array}[]{r}\lambda^{2}d^{T}Pd\leq\rho\\ (\lambda-1)^{2}d^{T}Pd\leq\rho\\ \end{array}\right.\Rightarrow\left\{\ \begin{array}[]{r}\lambda^{2}\|d\|_{P}^{2}\leq\rho\\ (\lambda-1)^{2}\|d\|_{P}^{2}\leq\rho\\ \end{array}\right.$ (19)
Solving for $\lambda$ we obtain
$\left\{\ \begin{array}[]{l}\lambda^{2}\leq\frac{c}{\|d\|_{P}^{2}}\\ \lambda^{2}\|d\|_{P}^{2}-2\lambda\|d\|_{P}^{2}+\|d\|_{P}^{2}\leq c\\ \end{array}\right.\Rightarrow 1-2\sqrt{\frac{c}{\|d\|_{P}^{2}}}\leq 0$ (20)
The two ellipsoids $\mathcal{E}_{c}(x_{sp,i})$ and $\mathcal{E}_{c}(x_{sp,j})$ intersect if
$2\sqrt{\frac{c}{\|d\|_{P}^{2}}}\geq 1\rightarrow\|d\|_{P}^{2}\leq 4c$ (21)
Remark 2
Although Algorithm 1 ensures safety according to Proposition 2, it does not guarantee target achievement. Agents within one-hop communication range may become immobilized at a certain position because any attempt to reach the target points could potentially violate the safety condition (11). We intentionally refrain from employing any enforcer to resolve this deadlock scenario, as our aim is to assess whether our MARL framework can autonomously generate target points that circumvent such situations.
In summary, our analysis provides the following key theoretical results:
  • Each agent’s setpoint update only impacts the safety of neighboring agents within its one-hop communication range. This localized effect is an important property that enables decentralized coordination.
  • we establish that Algorithm 1 guarantees the preservation of the safety condition in (11) for all agents,
  • We provide (21) a computationally tractable method for evaluating possible safety condition (11) violations.
Together, these theoretical results form the basis for our safe and distributed multi-agent control framework, ensuring that agents can dynamically update their setpoints while maintaining the prescribed safety guarantees through local communication and coordination.
Agent Type Safety Mechanism Explicit Penalty Truncation Avg Coverage Avg Safety Violations
Baseline A No No No 41% 13,880
Baseline B No Yes No 0.01% 0
Approach A Yes No No 39% 0
Approach B Yes Yes No 16% 0
Approach C Yes Yes Yes 22% 0
TABLE I: Comparison of agent performance, safety mechanisms, explicit penalties, and safety violations.
V Neural Network Architecture
We employ the neural network architecture for action-value estimation presented in [7], which is tailored to address the challenges of dynamic network bridging in a cooperative multi-agent setting, enabling the agents to extrapolate information about spatial and temporal dependencies while collaborating effectively. Such architecture combines Graph Neural Networks (GNNs) for relational modeling and exploits their message-passing mechanisms compatibly with the task optimized in the environment [Ref Section]. Furthermore, a Long Short-Term Memory (LSTM) is employed to cope with temporal dependencies and the partial observability of the task. The model is trained in a Centralized Training Decentralized Execution (CTDE) [15] fashion, where the agents optimize the same action-value function parameterization during training. However, during execution, each agent acts based on its local observations and the shared representations from its neighbors, adhering to the decentralized nature of the task.
We continue by describing the key components of the architecture below.
V-1 Graph Attention Layers
Graph Convolutional Reinforcement Learning (DGN) [16] facilitates cooperation between the agents by enabling them to share their latent representations with their immediate neighbors. To this end, we leverage Graph Attention Network (GAT) [17] to capture the spatial and relational dependencies between agents in the network. Specifically, we employ two Multi-Headed GAT that enable each agent to attend to and integrate information from its neighboring agents adaptively. This encoding allows the agents to condition their actions based on the dynamic network topology and inter-agent relationships effectively.
Additionally, we integrate target entities into the sharing process by equipping them with this same encoding module to produce their latent representation. If an agent’s neighborhood includes target entities, the agent can gather their representations and condition its actions accordingly.
V-2 Long Short-Term Memory
To handle the partial observability and temporal dynamics of the environment, we integrate a LSTM [18] layer into our architecture. The LSTM layer maintains a memory of past observations, allowing the agents to make informed decisions based on past interactions, partial observability, and the evolving environment.
During training, we employ observation stacking, aggregating observations over multiple time steps. Such a process gives the LSTM layer the necessary temporal context for effective learning in the presence of partial observability.
V-3 Dueling Action Decoder
To estimate action values effectively, we incorporate a Dueling Action Decoder [19] in our architecture. This component comprises separate streams for estimating state values and advantages, which are then combined to produce the final action value estimates.
VI Experimental Setup
We evaluated our approach in a simulated 2D environment with three agents and two moving targets. The environment was normalized with axes ranging from 0 to 1. Agents and targets had a communication range of 0.25, and their movement offset was set to 0.05 for the calculation of the next target point $w$. The physics simulator allowed us to simulate the non-linear Unmanned Aerial Vehicle (UAV) dynamics while we updated the positions of the agents in their paths toward the decided target point.
Our multi-agent strategies were trained across 1M training steps. Training episodes are truncated and Partial Episode Boostraping [20] was performed after 100 steps taken by each agent. Agents were initialized at fixed starting positions that satisfy safety condition (11), and the moving targets were randomly placed at a certain distance apart to ensure realistic difficulty and a feasible connection. Target motion and placing follow a seeded random pattern to control the scenarios during training and testing. The primary focus of our experiments was to demonstrate the effectiveness of our safe tracking control system in enabling the MARL approach to learn policies that satisfy safety constraints. To establish the importance of this contribution, we designed a set of experiments with varying training configurations:
  • Baseline A: We first trained a typical MARL approach without any safety tracking control, serving as a baseline.
  • Baseline B: To investigate the impact of explicit penalization for safety violations, we trained the same MARL approach as in Baseline A but with a high penalty (-100) imposed whenever an agent violated the safety constraints.
  • Approach A: This configuration represents our key contribution, where we trained the agents with the safe tracking control system enabled. Agents employed Algorithm 1 to update their target points while respecting the safety conditions.
The results from Approach A demonstrated the effectiveness of our safe tracking control system in enabling the agents to learn policies that accomplish the task while adhering to safety constraints.
To further solidify our approach and highlight its advantages, we conducted additional experiments:
  • Approach B: Here, we combined the safe tracking control system with explicit penalization for safety violations, similar to Baseline B. This configuration aimed to analyze the agents’ behavior when both safety measures and penalization for safety violations were employed simultaneously.
  • Approach C: Building upon Approach B, we additionally introduced episode termination whenever safety constraints were violated. This experiment assessed the impact of a more stringent safety enforcement mechanism.
If enabled, the safe tracking control system was used to prevent agents from continuing their movement if they violated the safety condition (11), such as going too close to other agents and risking collision. During training, in case the safety condition was violated, the agents involved were forced to stop their movement and would remain blocked until a new action toward safe directions was computed by at least one of the agents involved. To evaluate our agents, we ran 100 evaluation episodes disabling the safety tracking control system and counting the number of times that safety conditions were violated for every position update (14).
VII Results
Table I shows the results of our experiments evaluating different learned MARL strategies. We report the average communication coverage achieved by the agents as well as the number of safety violations observed when the safe tracking control system was disabled during evaluation.
Baseline A agents, trained without any safety mechanisms, achieved 41% coverage but incurred 13,880 safety violations on average. Introducing an explicit penalty of -100 for violating safety constraints (Baseline B) reduced the coverage to only 0.01% but did successfully prevent any safety violations.
In our Approach A, agents achieved 39% coverage (2% less than Baseline A) while completely avoiding any safety violations. This demonstrates that safe tracking control allows the agents to learn policies that respect safety constraints without significantly compromising task performance and without the need for explicit constraint violation penalties. Adding an explicit penalty for safety violations to the safe tracking control (Approach B) reduced coverage to 16% while still avoiding violations. Further adding episode truncation when violating safety (Approach C) increased coverage to 22% while maintaining zero violations.
VIII Discussion
Our work presents a novel hybrid approach that combines MARL with control-theoretic methods to address a complex cooperative task while ensuring safety guarantees. The theoretical analysis provides three key results. First, we prove that each agent’s setpoint update only affects the safety of its one-hop neighboring agents, enabling decentralized coordination. Second, we establish that Algorithm 1 guarantees the preservation of the safety condition for all agents under specific assumptions. Third, we derive an analytical condition to efficiently evaluate potential safety violations during setpoint updates.
The experimental results highlight the importance of our hybrid approach. Agents trained without any safety mechanisms (Baseline A) achieved high task coverage but incurred numerous safety violations. Introducing an explicit penalty for violating safety constraints (Baseline B) successfully prevented violations but at the cost of significantly compromised task performance. In contrast, our Approach A, which incorporates the safe tracking control system based on Algorithm 1, allowed agents to learn policies that respect safety constraints while maintaining reasonable task coverage, without explicit constraint violation penalties.
Despite these strengths, there are some limitations to our current work. While we have demonstrated the approach’s effectiveness on the network bridging task, we have not extensively evaluated its scalability to significantly larger swarm sizes or other domains. Moving forward we plan to investigate the scalability and adaptability of our approach across different domains and with varying swarm sizes.

### v2

@@
-Multi-Agent Reinforcement Learning with Control-Theoretic Safety Guarantees for Dynamic Network Bridging
+Safety-Aware Multi-Agent Learning for Dynamic Network Bridging
 Abstract
 Addressing complex cooperative tasks in safety-critical environments poses significant challenges for multi-agent systems, especially under conditions of partial observability.
-This work introduces a hybrid approach that integrates Multi-Agent Reinforcement Learning with control-theoretic methods to ensure safe and efficient distributed strategies.
-Our contributions include a novel setpoint update algorithm that dynamically adjusts agents’ positions to preserve safety conditions without compromising the mission’s objectives.
-Through experimental validation, we demonstrate significant advantages over conventional MARL strategies, achieving comparable task performance with zero safety violations.
-Our findings indicate that integrating safe control with learning approaches not only enhances safety compliance but also achieves good performance in mission objectives.
+We focus on a dynamic network bridging task, where agents must learn to maintain a communication path between two moving targets.
+To ensure safety during training and deployment, we integrate a control-theoretic safety filter that enforces collision avoidance through local setpoint updates.
+We develop and evaluate multi-agent reinforcement learning safety-informed message passing, showing that encoding safety filter activations as edge-level features improves coordination.
+The results suggest that local safety enforcement and decentralized learning can be effectively combined in distributed multi-agent tasks.
 I Introduction
@@
 Moreover, the inherent nature of multi-agent systems (MAS) often involves partial observability, further increasing the complexity of these problems by limiting the information available to each agent about the environment and the states of other agents.
-This challenge highlights a fundamental limitation of all Reinfocement Learning (RL)-based approaches: while agents learn to optimize predefined reward functions, their behavior lacks guarantees and remains inherently unpredictable.
+This challenge highlights a fundamental limitation of all reinforcement learning (RL)-based approaches: while agents learn to optimize predefined reward functions, their behavior lacks guarantees and remains inherently unpredictable.
 The sole reliance on MARL’s reward functions to ensure safety is insufficient [1].
@@
 In ensuring safety among agents, various control-theoretic approaches and hybrid learning approaches have been developed, ranging from Model Predictive Control (MPC) [2] to Reference Governor (RG) techniques [3], and from Control Barrier Functions (CBFs) [4] to Reachability Analysis (RA) methods [5].
-In this work, we focus on developing a hybrid learning approach built upon the method used in [6], that can learn to achieve complex objectives while ensuring safety guarantees in distributed MAS.
-To demonstrate and evaluate the effectiveness of our hybrid approach we consider the cooperative task of Dynamic Network Bridging [7] (Fig.
-1).
-In this task the goal is to establish and maintain a connection between two moving targets A and B in a 2D plane, relying on a swarm of $N$ agents with limited observation of the environment.
-The agents must form an ad-hoc mobile network that establishes a communication path between the two targets as they move, dynamically adjusting their positions to ensure an uninterrupted multi-hop communication link.
+In robotics applications, those methods are known as safety filters [6], and they are deployed to ensure safety regardless of the algorithm used for the agent to accomplish a task.
+While safety filters can effectively prevent unsafe behavior, they are often conservative by design.
+This conservatism may lead to conflicts with task objectives, especially in MAS where coordination is essential.
+In this work, we develop a training framework for MARL that incorporates safety constraints during policy learning.
+We adapt an optimization-free safety tracking control algorithm developed in [7], which is based on ellipsoidal positively invariant sets.
+We show how this safety filter can be adapted to the context of MAS and how safety is guaranteed.
+We integrate this type of filter in the agents’ training process, allowing them to learn decentralized policies and avoid situations where the filter compromises task objectives.
 Figure 1: Decentralized swarm coordination for dynamic network bridging
-In this paper we introduce a hybrid approach combining MARL and control-theoretic methods, accomplishing task objectives while providing safety guarantees.
-Our key contributions are: 1) A decentralized control framework that is MARL compatible and restricts the effect of each agent’s movement updates to only its one-hop neighbors, enabling efficient local coordination; 2) An algorithm that updates setpoints while preserving safety conditions through communication with affected neighbors; 3) An analytical computationally tractable condition verifying potential safety violations during updates.
+We demonstrate our approach in the dynamic network bridging task [8] (Fig.
+1) where a number of agents establish and maintain a communication path between two moving targets.
+Our key contributions are as follows: 1) A novel decentralized control framework that integrates a safety filter based on invariant sets into the MARL agents’ training process, and restricts the effect of each agent’s movement updates to only its one-hop neighbors, enabling efficient local coordination; 2) We augment the learning process with safety-informed message passing, where agents observe safety filter activations as edge-level features during training, leading to improved coordination 3) An algorithm that updates setpoints while preserving safety conditions through communication with affected neighbors; 3) An analytical computationally tractable condition verifying potential safety violations during updates.
 Overall our approach provides effective coordination among agents while ensuring safety constraints.
 II Preliminaries
-In this section, we present the foundational theoretical elements, drawing from both control theory and MARL, that form the basis of our proposed hybrid approach.
-Consider a linear time invariant (LTI) system expressed in the state space form
-$\dot{x}=Ax+Bu$ (1)
-where $x\in\mathbb{R}^{n}$ is the state vector with $n$ components, and $u\in\mathbb{R}^{p}$ is the input vector with $p$ components.
-Thus, the dynamical matrices are $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times p}$.
-If the pair $(A,B)$ is controllable, there exists a state feedback controller:
-$u=-Kx$ (2)
-such that, the origin ($x=0$) of the resulting closed-loop (C-L) system:
-$\dot{x}=(A-BK)x=A_{f}x$ (3)
-is an asymptotically stable equilibrium point (EP): $x(t)\rightarrow 0\quad\text{as}\quad t\rightarrow+\infty$.
-Hence, the matrix $A_{f}$ is Hurwitz, i.e.
-all the eigenvalues have a negative real part [8].
-The problem to design (2) that makes (1) asymptotically stable is called stabilization problem.
-II-A Tracking Control Problem
-In the case the problem requires to move the system (1) following a specific trajectory, the problem is called tracking control.
-This problem can be solved by generating a sequence of EPs [9] that we call setpoints, denoted by $x_{sp}$.
-The C-L system (3) can be represented as:
-$\dot{x}=A_{f}(x-x_{sp})$ (4)
-where the control input (2) becomes $u=-K(x-x_{sp}).$
-II-B Lyapunov’s Theory
-Since $A_{f}$ is Hurwitz, there exists a symmetric positive definite matrix $P>0$ such that:
-$A_{f}^{T}P+PA_{f}=-Q$ (5)
-for a given symmetric positive definite matrix $Q>0$. Equation (5) is called Lyapunov’s equation. Defining the $P$-norm as:
-$\lVert x-x_{sp}\rVert_{P}\triangleq\sqrt{(x-x_{sp})^{T}P(x-x_{sp})}$ (6)
-the $P$-norm squared is a Lyapunov function $V(x)=\lVert x-x_{sp}\rVert_{P}^{2}$, which is always positive except at $x_{sp}$ where it is zero.
-For all trajectories $x(t)$ starting within the ellipsoid:
-$\mathcal{E}_{c}(x_{sp})=\{x\in\mathbb{R}^{n}\,|\,\lVert x-x_{sp}\rVert_{P}^{2}\leq c\},$ (7)
-where $c>0$, $V(x(t))$ is monotonically decreasing. Therefore, the trajectory $x(t)$ never leaves $\mathcal{E}_{c}$, which is called a positively invariant set [8].
-Remark 1
-In the case of a nonlinear dynamical system described by $\dot{x}=f(x,u)$ where $u=0$, $x=0$ is an equilibrium point for the system $f(0,0)=0$.
-The control input (2) can be designed by considering the linearized system (1) around the equilibrium point.
-Therefore, the closed-loop (CL) system (3) can be used to solve Lyapunov’s equation (5), which is used to compute the positively invariant set $\mathcal{E}_{c}(x_{sp})$ (7).
-While for LTI systems, $\mathcal{E}_{c}(x_{sp})$ is positively invariant for any positive $c\in\mathbb{R}$, for nonlinear systems, we need to verify that condition in a neighborhood of the equilibrium point $x_{sp}$ [8].
-Furthermore, for nonlinear systems, $x_{sp}$ has to satisfy $f(x_{sp},0)=0$.
-II-C 6-DOF UAV
-In this work, we consider a 6-DOF UAV as the agent, described by a state vector of 12 components:
-$x=\begin{bmatrix}\dot{p}_{x},&p_{x},&\dot{p}_{y},&p_{y},&\dot{p}_{z},&p_{z},&\dot{\phi},&\phi,&\dot{\theta},\\ \theta,&\dot{\psi},&\psi\end{bmatrix}^{T}$ (8)
-where $p_{x}$, $p_{y}$, $p_{z}$, and their derivatives represent linear position and velocity, respectively.
-Similarly, $\phi$, $\theta$, $\psi$, and their derivatives represent the Euler angles (roll, pitch, and yaw) and their angular velocities, respectively.
-This model is consistent with the one used in [9] and references therein.
-The controller used is given by (2), obtained by linearizing the quadrotor model around an equilibrium point.
-For this case, the equilibrium point is indicated with $x_{sp}$ which has the same structure of (8) with all components equal to zero except for $p_{x}$, $p_{y}$, $p_{z}$.
-II-D Multi-Agent Reinforcement Learning
+II-A Task description
+Dynamic network bridging is a multi-agent task focused on addressing the problem of dynamically establishing and maintaining a communication link between two mobile, independent nodes (targets) using a swarm of autonomous agents taking decentralized decisions [8].
+Agents and targets are moving in a 2D grid space and have a limited communication range, within which they exchange information.
+This region can be defined as a 2D ball of radius $r$, centered at its position $p_{i}(t)=[p_{x,i}(t),p_{y,i}(t),p_{z,i}(t)]^{T}$.
+Based on this definition, two agents $i$ and $j$ can communicate if and only if their regions intersect:
+$\mathcal{B}_{r}(p_{i}(t))\cap\mathcal{B}_{r}(p_{j}(t))\neq\emptyset.$ (1)
+where
+$\mathcal{B}_{r}(p_{i}(t))=\left\{p\in\mathbb{R}^{3}:\|p-p_{i}(t)\|^{2}\leq r^{2}\right\}$
+In our setup, we consider $p_{z,i}$ fixed, therefore the one-hop communication region can be described by a 2D ball.
+Formally, the system is represented as a dynamic graph $\mathcal{G}(t)=(\mathcal{V},\mathcal{E}(t))$ with nodes $\mathcal{V}$ corresponding to agents and targets.
+An edge $(u,v)$ exists between nodes if their Euclidean distance at time-step $t$ is within a predefined communication range.
+The task reflects scenarios where communication infrastructure is absent or unreliable, such as in disaster zones or remote environments.
+Agents must position themselves to form an ad-hoc, decentralized communication network that connects the two mobile targets.
+Each agent operates based on local observations and limited-range communication, with no access to global state information.
+II-B Multi-agent reinforcement learning
 In a typical multi-agent cooperative scenario, $N$ agents interact within an environment to complete a task.
@@
 A Dec-POMDP is defined by a tuple $(S,\mathbf{A},\mathcal{O},T,R,\Omega,\gamma,N)$ where $S$ is a finite set of states of the environment, $\mathbf{A}$ is a set of joint action spaces, and $N$ is the number of agents, $T$ is the state transition probability function which gives the probability of transitioning from state $s$ to state $s^{\prime}$ when joint action $\mathbf{a}=(a_{1},\cdots,a_{N})$ is taken; $R$ is the joint reward function that maps states and joint actions to real numbers and is used to specify the goal of the agents; $\Omega$ is a set of joint observations $O_{i}$, where $O_{i}$ is the observation set for agent $i$; $\mathcal{O}$ is the joint observation probability function, which gives the probability of receiving joint observation $\mathbf{o}=(o_{1},\cdots,o_{N})$ after taking joint action $\mathbf{a}$ and ending up in state $s^{\prime}$; and $\gamma$ is the reward discount factor, $0\leq\gamma\leq 1$.
-In a Dec-POMDP, agents need to collaborate and coordinate their actions based on their individual observations to maximize the joint reward function.
-Solving a Dec-POMDP optimally is computationally intractable, as the policy space grows exponentially with the number of agents and the planning horizon.
-Therefore, various approximate solution methods [12, 13] have been used, such as heuristic search, dynamic programming, value or policy-based approximators with neural networks etc.
-The problem facing the team is to find the optimal joint policy, i.e.
-a combination of individual agent policies that produces behavior that maximizes the team’s expected reward.
-III Problem Formulation
-We address the challenge of dynamically forming and sustaining a connection between two moving targets, A and B, in a 2D space using a swarm of $N$ agents [7].
-These agents must collaboratively adjust their positions to maintain a continuous multi-hop link between the targets, operating under decentralized control and limited local perception, while ensuring they avoid collisions and respect physical boundaries.
-We consider all the agents as nonlinear systems that can be reduced into the form (1) controlled by (2) around the equilibrium point $x_{sp,i}$ [9]:
-$\dot{x}_{i}=A_{f}(x_{i}-x_{sp,i})=(A-BK)(x_{i}-x_{sp,i})$ (9)
-for $i=1,\ldots,N$.
-In this case, we consider $N$ agents described with the same model, but in general, our approach can be generalized to the case of different dynamics and multidimensional scenarios.
-III-A One-hop Communication
-We assume that for each agent, there is a ball of radius $r$ in the 2- or 3-dimensional Euclidean space centered at the agent’s current linear position $p_{i}(t)=[p_{x,i}(t),p_{y,i}(t),p_{z,i}(t)]^{T}$:
-$\mathcal{B}_{r}(p_{i}(t))=\left\{p\in\mathbb{R}^{3}:\|p-p_{i}(t)\|^{2}\leq r^{2}\right\}$
-Given two agents $i$ and $j$ with $i\neq j$, they can communicate if and only if
-$\mathcal{B}_{r}(p_{i}(t))\cap\mathcal{B}_{r}(p_{j}(t))\neq\emptyset$ (10)
-In our setup, since $p_{z,i}$ is fixed, the one-hop communication region is described by a 2-D ball.
-Note that, we use a different notation when we refer to the 2-D communication space to avoid any confusion between the quadrotor $x$-axis and the state space $x$.
-III-B Safety
+The goal of the agents is to maximize a joint reward function.
+Solving a Dec-POMDP optimally is computationally intractable, therefore, various approximate solution methods [11, 12] have been used, such as heuristic search, dynamic programming, value or policy-based approximators with neural networks etc.
+II-C Distributed autonomous swarm formation for dynamic network bridging
+We formulate our problem as a Dec-POMDP, where each agent has restricted sensing capabilities given by a specific communication range, leading to incomplete information about the global state of the system outside such area.
+However, whenever two entities (agents or targets) are within range, a communication channel becomes available for them to exchange information.
+This channel enables the possibility for agents to know each other features and execute distributed strategies requiring message exchange to coordinate.
+At the beginning of every time-step, agents make decentralized decisions, generating target points $w$, utilizing their local observation to choose the direction of their next movement.
+To solve the Dec-POMDP that corresponds to our task we used a graph neural network (GNN) approximator that encodes the graph structure of the problem.
+II-D Agent Dynamics
+We consider a linear time-invariant (LTI) of the form
+$\dot{x}=A(x-x_{sp})$ (2)
+where $x\in\mathbb{R}^{n}$ is the state of the system, and the origin is translated to the setpoint $x_{sp}\in\mathbb{R}^{n}$.
+The matrix $A\in\mathbb{R}^{n\times n}$ is Hurwitz (i.e., $\mathrm{Re}(\lambda_{i})<0$), where $\lambda_{i}\in\mathbb{C}$ is the $i$-th eigenvalue of $A$.
+The matrix $A$ can be seen as the closed-loop system matrix obtained by applying a state-feedback controller (i.e., $u=-K(x-x_{sp})$), where $u$ is the control input.
+For the asymptotically stable LTI system (2), there exists a quadratic Lyapunov function of the form
+$V(x)=(x-x_{sp})^{T}P(x-x_{sp})$ (3)
+where $P$ is a symmetric positive definite matrix that satisfies the Lyapunov equation
+$A^{T}P+PA=-Q$ (4)
+for a given symmetric positive definite matrix $Q>0$.
+Defining the $P$-norm as
+$\lVert x-x_{sp}\rVert_{P}\triangleq\sqrt{(x-x_{sp})^{T}P(x-x_{sp})},$ (5)
+for a fixed value $c>0$, the Lyapunov function (3) defines an ellipsoid, which can be written in terms of (5) as follows:
+$\mathcal{E}_{c}(x_{sp})=\{x\in\mathbb{R}^{n}\mid\lVert x-x_{sp}\rVert_{P}^{2}\leq c\}.$ (6)
+From (4), $V(x(t))$ is a monotonically decreasing function; therefore, for any $x(t)$ starting within the ellipsoid (6), the trajectory never leaves it.
+Hence, $\mathcal{E}_{c}$ is called a positively invariant set [13].
+For the rest of the paper, we will refer to the state of the agent $i$ as $x_{i}$ and for the related setpoint as $x_{sp,i}$.
+In this context, the tracking algorithm updates the setpoint of the agent to reach a desired position.
+This formulation can be extended to more general nonlinear systems by using linearization [13], or Lyapunov-like methods such as quadratic boundedness [14], as used in the context of UAVs in [7], where ellipsoidal positively invariant sets are provided.
+We use this last case in our experiments.
+II-E Safety
 Definition 1
-Given two agents $i$ and $j$, where $x_{sp,i}\neq x_{sp,j}$, $x_{i}(t_{0})\in\mathcal{E}_{c}(x_{sp,i})$, and $x_{j}(t_{0})\in\mathcal{E}_{c}(x_{sp,j})$.
-Agents $i$ and $j$ are safe with respect to each other if
+Given two agents $i$ and $j$, with distinct equilibrium points $x_{sp,i}\neq x_{sp,j}$ and initial states $x_{i}(t_{0})\in\mathcal{E}_{c}(x_{sp,i})$ and $x_{j}(t_{0})\in\mathcal{E}_{c}(x_{sp,j})$, respectively, agents $i$ and $j$ are deemed safe relative to each other if:
 $\mathcal{E}_{c}(x_{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})=\emptyset.$ (7)
-The above definition is derived by the positively invariant property of (7), since the initial conditions of each agent are inside their respective ellipsoids, then $x_{i}(t)\in\mathcal{E}_{c}(x_{sp,i})$ and $x_{j}(t)\in\mathcal{E}_{c}(x_{sp,j})$ for $t\rightarrow+\infty$.
-This means that the trajectory of each agent cannot intersect generating the risk of possible collisions.
-In order to guarantee safety in our setup, we need the following assumption:
+This definition leverages the invariant properties of (6), ensuring agents’ trajectories remain confined within their respective safety ellipsoids, thereby preventing collisions.
+To operationalize this safety criterion within the two-dimensional communication space, we impose the following:
 Assumption 1
-The projection of $\mathcal{E}_{c}(x_{sp,i})$ into the 2-D space formed by the components $p_{x}$ and $p_{y}$ must be contained within the 2-dimensional ball $\mathcal{B}_{r}(p_{i}(t))$, where $p_{i}(t)=[p_{x,i}(t),p_{y,i}(t)]$ represents the actual position coordinates ($p_{x,i}(t),p_{y,i}(t)$) of agent $i$ for any state $x_{i}\in\mathcal{E}_{c}(x_{sp,i})$.
-III-C Cooperative MARL for Dynamic Network Bridging
-Each agent $i$ has a local observation $o_{i}$ comprising the structure of its neighborhood and features representing its neighbors at certain time-step $t$.
-Such features include the node ID, current coordinates, action taken, and coordinates of the targets A and B to be connected.
-At the beginning of every time-step, agents make decentralized decisions, generating target points $w$, utilizing their local observation to choose the direction of their next movement.
-This is encoded in a two-dimensional action space, with each dimension having three options: move forward, move backward, or hold along the corresponding $p_{x}$ or $p_{y}$ axis.
-Given the direction chosen by the agents, a new target point is calculated for each agent using a fixed offset equal for both the $p_{x}$ and $p_{y}$ axes.
-Finally, agents are rewarded using a reward function that motivates agents to form a connected network bridging the targets.
-Such reward signal combines three components:
-Base Connectivity Reward: The base reward, $R_{\text{base}}$, encourages larger connected components:
-$R_{\text{base}}(s)=\frac{|C_{\text{max}}(s)|}{|\mathcal{V}|},$ (12)
-where $|C_{\text{max}}(s)|$ represents the size of the largest connected component in state $s$, and $|\mathcal{V}|$ is the total number of entities.
-Centroid Distance Penalty: The centroid distance penalty, $P_{\text{cent}}$, is based on agents’ positions relative to the targets, penalizing them based on the Euclidean distance between the centroid of the agents’ positions and the central point between the targets.
-Target Path Bonus: A bonus, $B_{\text{path}}=100$, is awarded if a path exists between the two targets.
-The overall reward combines these three elements:
-$R(s,a)=\begin{cases}B_{\text{path}}(s)&\mbox{if }\exists\mathrm{path}(T_{1},T_{2});\par\\ R_{\text{base}}(s)-P_{\text{cent}}(s),&\mbox{otherwise}.\end{cases}$ (13)
-We also include a physics simulator, modeling the continuous non-linear dynamics of the agents (2), and a safe tracking control system proposed in the next section (Algorithm 1).
-IV Safe Tracking Control
-The safety tracking control algorithm that we propose here consists of two steps, the first one is the setpoint update of the agent $i$ that guarantees that its current position belongs to $\mathcal{E}_{c}$ of the new setpoint.
-This condition is fundamental for the second step where the safety condition (11) is checked for all agents within the one-hop communication ball.
+The projection of $\mathcal{E}_{c}(x_{sp,i})$ onto the 2-D plane defined by the coordinates $(p_{x},p_{y})$ must always lie within the 2-D communication ball $\mathcal{B}_{r}(p_{i}(t))$, where $p_{i}(t)=[p_{x,i}(t),p_{y,i}(t)]$ represents the actual position of agent $i$.
+III Safety Filter for Tracking Control
+The proposed algorithm consists of two steps: first, computing a feasible safe step (Figure 2(a)); and second, checking whether the new position would interfere with another agent (Figure 2(b)).
 Let us start with the first step which has been borrowed from [7].
@@
 We can say that the agent has reached the setpoint $x_{sp,i}$ if and only if $x_{i}(t)\in\mathcal{E}_{s}(x_{sp,i})$.
-Assuming that the agent $i$ is moving from the target point $w_{i}$ to the target point $w_{i+1}$, generated by the MARL, the idea is to generate a sequence of setpoints on the line that connects the two target points.
+Assuming that the agent $i$ is moving from the target point $w_{i}$ to the target point $w_{i+1}$, generated by its decentralized policy, the idea is to generate a sequence of setpoints on the line that connects the two target points.
 Once the state of agent $i$, $x_{i}$ reached the setpoint $x_{sp,i}$, a new setpoint $x^{\prime}_{sp,i}$ is generated following the rule
@@
 Fig.
-2 illustrates the process of updating the setpoint, presented in 2-D space for clearer visualization, though it actually corresponds to a state space with n=12 components, as applicable to quadrotors.
-Figure 2: Setpoint update for agent $i$ in the state space. Note that the target points $w_{i}$ and $w_{i+1}$ have the same structure of the setpoints $x_{sp,i}$.
+2(a) illustrates the process of updating the setpoint, presented in 2-D space for clearer visualization, though it actually corresponds to a state space with n=12 components, as applicable to quadrotors.
+(a) Setpoint update for agent $i$ in the state space.
+(b) $m_{\lambda}$ is on the segment that connects the two setpoints.
+Figure 2: Safe tracking control and safety verification.
 Thanks to (8) and Assumption 1, the projection of $\mathcal{E}_{c}(x^{\prime}_{sp,i})$ is contained in $\mathcal{B}(p_{i}(t))$.
@@
 Proof:
-From Assumption 1, the ball $\mathcal{B}_{r}(p_{i}(t))$ encompasses the projection onto the 2-D space of $\mathcal{E}_{c}(x{sp,i})$ for any current state $x_{i}(t)\in\mathcal{E}_{c}(x_{sp,i})$.
+From Assumption 1, the ball $\mathcal{B}_{r}(p_{i}(t))$ encompasses the projection onto the 2-D space of $\mathcal{E}_{c}(x_{sp,i})$ for any current state $x_{i}(t)\in\mathcal{E}_{c}(x_{sp,i})$.
 Utilizing (8) to adjust the setpoint ensures that $x_{i}(t)$ remains within $\mathcal{E}_{c}(x^{\prime}_{sp,i})$, signifying that its 2-D space projection still falls within $\mathcal{B}_{r}(p_{i}(t))$.
@@
 ∎
-Proposition 1 highlights a critical aspect of setpoint updates: setpoint updates only impact safety for agents within one-hop range.
-This ensures agent $i$ can communicate with affected neighbors before executing updates.
+Proposition 1 highlights a critical aspect of setpoint updates: setpoint updates only impact safety for agents within communication range.
+This ensures agent $i$ can exchange messages with affected neighbors before executing updates.
 This allows coordinating to preemptively address any potential safety concerns from the new setpoint through direct communication with impacted agents.
@@
 Output: Updated setpoint $x_{sp,i}$
-safety_violation $\leftarrow$ 0;
+potential_safety_violation $\leftarrow$ 0;
 if $x_{i}(t)\in\mathcal{E}_{s}(x_{sp,i})$ then
@@
              if $\mathcal{E}_{c}(x_{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})\neq\emptyset$  then
-                   safety_violation $\leftarrow 1$;
+                   potential_safety_violation $\leftarrow 1$;
                    break;
@@
        end for
-       if safety_violation ==absent$=$ = 00 then
+       if potential_safety_violation ==absent$=$ = 00 then
              $x_{sp,i}\leftarrow x^{\prime}_{sp,i}$;
@@
 Proposition 2
-Considering the multi-agent system described by (9) and Assumptions 1, 2 and 3, then applying Algorithm 1 to each agent guarantees that the safety condition (11) is satisfied for all agents.
+Considering the MAS described where the agents are LTI systems (2) and Assumptions 1, 2 and 3, then applying Algorithm 1 to each agent guarantees that the safety condition (7) is satisfied for all agents.
 Proof:
@@
 ∎
-We now need to verify when the safety condition (11) is violated.
+In the second step, we need to verify whether the safety condition (7) is going to be violated.
 Our objective is to demonstrate that if (7) is violated, then there exists a point on the line connecting $x_{sp,i}$ and $x_{sp,j}$, inside the overlap region of two ellipsoids $\mathcal{E}_{c}(x_{sp,i})$ and $\mathcal{E}_{c}(x_{sp,j})$ (Fig.
-3).
+2(b)).
 This overlap region, denoted by $\mathcal{E}_{c}(x{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})$, satisfies the following condition:
@@
 See [15]. ∎
-Proposition 3 shows that if there is an ellipsoid violation, then the point $m_{\lambda}$ belongs to the intersection Fig. 3:
+Proposition 3 shows that if there is an ellipsoid violation, then the point $m_{\lambda}$ belongs to the intersection Fig. 2(b):
 $\mathcal{E}_{c}(x_{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})\neq\emptyset\rightarrow m_{\lambda}\in\mathcal{E}_{c}(x_{sp,i})\cap\mathcal{E}_{c}(x_{sp,j})$
-Figure 3: Explaination of Proposition 3: $m_{\lambda}$ is on the segment that connects the two setpoints.
 Therefore, to test if the two ellipsoids intersect we need to find if there exists at least one point on the segment joining $x_{sp,i}$ and $x_{sp,j}$ that satisfies (10).
@@
 Agents within one-hop communication range may become immobilized at a certain position because any attempt to reach the target points could potentially violate the safety condition (7).
-We intentionally refrain from employing any enforcer to resolve this deadlock scenario, as our aim is to assess whether our MARL framework can autonomously generate target points that circumvent such situations.
+We intentionally refrain from employing any enforcer to resolve deadlock scenarios.
+Instead, we allow the agents to learn how to avoid or recover from them through the training process.
 In summary, our analysis provides the following key theoretical results:
@@
 Together, these theoretical results form the basis for our safe and distributed multi-agent control framework, ensuring that agents can dynamically update their setpoints while maintaining the prescribed safety guarantees through local communication and coordination.
-Agent Type Safety Mechanism Explicit Penalty Truncation Avg Coverage Avg Safety Violations
-Baseline A No No No 41% 13,880
-Baseline B No Yes No 0.01% 0
-Approach A Yes No No 39% 0
-Approach B Yes Yes No 16% 0
-Approach C Yes Yes Yes 22% 0
-TABLE I: Comparison of agent performance, safety mechanisms, explicit penalties, and safety violations.
-V Neural Network Architecture
-We employ the neural network architecture for action-value estimation presented in [7], which is tailored to address the challenges of dynamic network bridging in a cooperative multi-agent setting, enabling the agents to extrapolate information about spatial and temporal dependencies while collaborating effectively.
-Such architecture combines Graph Neural Networks (GNNs) for relational modeling and exploits their message-passing mechanisms compatibly with the task optimized in the environment [Ref Section].
-Furthermore, a Long Short-Term Memory (LSTM) is employed to cope with temporal dependencies and the partial observability of the task.
-The model is trained in a Centralized Training Decentralized Execution (CTDE) [15] fashion, where the agents optimize the same action-value function parameterization during training.
-However, during execution, each agent acts based on its local observations and the shared representations from its neighbors, adhering to the decentralized nature of the task.
-We continue by describing the key components of the architecture below.
-V-1 Graph Attention Layers
-Graph Convolutional Reinforcement Learning (DGN) [16] facilitates cooperation between the agents by enabling them to share their latent representations with their immediate neighbors.
-To this end, we leverage Graph Attention Network (GAT) [17] to capture the spatial and relational dependencies between agents in the network.
-Specifically, we employ two Multi-Headed GAT that enable each agent to attend to and integrate information from its neighboring agents adaptively.
-This encoding allows the agents to condition their actions based on the dynamic network topology and inter-agent relationships effectively.
-Additionally, we integrate target entities into the sharing process by equipping them with this same encoding module to produce their latent representation.
-If an agent’s neighborhood includes target entities, the agent can gather their representations and condition its actions accordingly.
-V-2 Long Short-Term Memory
-To handle the partial observability and temporal dynamics of the environment, we integrate a LSTM [18] layer into our architecture.
-The LSTM layer maintains a memory of past observations, allowing the agents to make informed decisions based on past interactions, partial observability, and the evolving environment.
-During training, we employ observation stacking, aggregating observations over multiple time steps.
-Such a process gives the LSTM layer the necessary temporal context for effective learning in the presence of partial observability.
-V-3 Dueling Action Decoder
-To estimate action values effectively, we incorporate a Dueling Action Decoder [19] in our architecture.
-This component comprises separate streams for estimating state values and advantages, which are then combined to produce the final action value estimates.
-VI Experimental Setup
-We evaluated our approach in a simulated 2D environment with three agents and two moving targets.
-The environment was normalized with axes ranging from 0 to 1.
-Agents and targets had a communication range of 0.25, and their movement offset was set to 0.05 for the calculation of the next target point $w$.
-The physics simulator allowed us to simulate the non-linear Unmanned Aerial Vehicle (UAV) dynamics while we updated the positions of the agents in their paths toward the decided target point.
-Our multi-agent strategies were trained across 1M training steps.
-Training episodes are truncated and Partial Episode Boostraping [20] was performed after 100 steps taken by each agent.
-Agents were initialized at fixed starting positions that satisfy safety condition (11), and the moving targets were randomly placed at a certain distance apart to ensure realistic difficulty and a feasible connection.
-Target motion and placing follow a seeded random pattern to control the scenarios during training and testing.
-The primary focus of our experiments was to demonstrate the effectiveness of our safe tracking control system in enabling the MARL approach to learn policies that satisfy safety constraints.
-To establish the importance of this contribution, we designed a set of experiments with varying training configurations:
-  • Baseline A: We first trained a typical MARL approach without any safety tracking control, serving as a baseline.
-  • Baseline B: To investigate the impact of explicit penalization for safety violations, we trained the same MARL approach as in Baseline A but with a high penalty (-100) imposed whenever an agent violated the safety constraints.
-  • Approach A: This configuration represents our key contribution, where we trained the agents with the safe tracking control system enabled.
-Agents employed Algorithm 1 to update their target points while respecting the safety conditions.
-The results from Approach A demonstrated the effectiveness of our safe tracking control system in enabling the agents to learn policies that accomplish the task while adhering to safety constraints.
-To further solidify our approach and highlight its advantages, we conducted additional experiments:
-  • Approach B: Here, we combined the safe tracking control system with explicit penalization for safety violations, similar to Baseline B.
-This configuration aimed to analyze the agents’ behavior when both safety measures and penalization for safety violations were employed simultaneously.
-  • Approach C: Building upon Approach B, we additionally introduced episode termination whenever safety constraints were violated.
-This experiment assessed the impact of a more stringent safety enforcement mechanism.
-If enabled, the safe tracking control system was used to prevent agents from continuing their movement if they violated the safety condition (11), such as going too close to other agents and risking collision.
-During training, in case the safety condition was violated, the agents involved were forced to stop their movement and would remain blocked until a new action toward safe directions was computed by at least one of the agents involved.
-To evaluate our agents, we ran 100 evaluation episodes disabling the safety tracking control system and counting the number of times that safety conditions were violated for every position update (14).
-VII Results
-Table I shows the results of our experiments evaluating different learned MARL strategies.
-We report the average communication coverage achieved by the agents as well as the number of safety violations observed when the safe tracking control system was disabled during evaluation.
-Baseline A agents, trained without any safety mechanisms, achieved 41% coverage but incurred 13,880 safety violations on average.
-Introducing an explicit penalty of -100 for violating safety constraints (Baseline B) reduced the coverage to only 0.01% but did successfully prevent any safety violations.
-In our Approach A, agents achieved 39% coverage (2% less than Baseline A) while completely avoiding any safety violations.
-This demonstrates that safe tracking control allows the agents to learn policies that respect safety constraints without significantly compromising task performance and without the need for explicit constraint violation penalties.
-Adding an explicit penalty for safety violations to the safe tracking control (Approach B) reduced coverage to 16% while still avoiding violations.
-Further adding episode truncation when violating safety (Approach C) increased coverage to 22% while maintaining zero violations.
-VIII Discussion
-Our work presents a novel hybrid approach that combines MARL with control-theoretic methods to address a complex cooperative task while ensuring safety guarantees.
-The theoretical analysis provides three key results.
-First, we prove that each agent’s setpoint update only affects the safety of its one-hop neighboring agents, enabling decentralized coordination.
-Second, we establish that Algorithm 1 guarantees the preservation of the safety condition for all agents under specific assumptions.
-Third, we derive an analytical condition to efficiently evaluate potential safety violations during setpoint updates.
-The experimental results highlight the importance of our hybrid approach.
-Agents trained without any safety mechanisms (Baseline A) achieved high task coverage but incurred numerous safety violations.
-Introducing an explicit penalty for violating safety constraints (Baseline B) successfully prevented violations but at the cost of significantly compromised task performance.
-In contrast, our Approach A, which incorporates the safe tracking control system based on Algorithm 1, allowed agents to learn policies that respect safety constraints while maintaining reasonable task coverage, without explicit constraint violation penalties.
-Despite these strengths, there are some limitations to our current work.
-While we have demonstrated the approach’s effectiveness on the network bridging task, we have not extensively evaluated its scalability to significantly larger swarm sizes or other domains.
-Moving forward we plan to investigate the scalability and adaptability of our approach across different domains and with varying swarm sizes.
+Training Conditions Evaluation Results
+Safety Filter Explicit Penalty Truncation Avg Coverage Avg Safety Interventions
+Baseline No No No $(37.63\pm 16.62)\%$ $26.99\pm 12.31$
+Approach A Yes No No $(50.14\pm 12.72)\%$ $20.45\pm 7.49$
+Approach B Yes Yes No $(45.02\pm 11.81)\%$ $15.25\pm 7.82$
+Approach C Yes Yes Yes $(10.86\pm 6.44)\%$ $9.76\pm 5.72$
+TABLE I: Comparison of the trained agents performance across 100 evaluation episodes.
+IV Decentralized Safety-Aware Agents
+This section details the components of our approach to ensure safe interactions while learning effective positioning strategies in networked and decentralized multi-agent tasks, by combining MARL with our safety filter mechanism.
+The core intuition is to employ Algorithm 1 since the training phase and inform the agents of eventual safety filter activations.
+IV-A Guiding learning via decentralized safety filters
+By enforcing adherence to explicit safety conditions defined in (7), agents experience realistic operational constraints and encounter possible deadlock scenarios whenever direct trajectories toward their target points trigger the safety filter.
+These constraints inherently shape the policy optimization landscape during training: 1) creating deadlock conditions agents might be blocked by their respective safety filters unless they reconsider their actions; 2) avoiding unsafe states of the environment, which are not going to be visited by the agents.
+Consequently, the agents need to leverage interactions and communication with neighboring agents to collaboratively avoid or resolve deadlock situations by generating alternative, safety-compliant paths to achieve their objectives effectively.
+IV-B Safety-informed edges
+Leveraging the graph nature of the task introduced in Section II-C, we explicitly introduce edge-level features linked to the conditions determining the activation of the decentralized safety filter.
+Specifically, edge features comprise base attributes controlling the activation of the safety filter and a binary safety activation indicator between connected agents.
+Formally, given two connected agents $i$ and $j$, the edge features are defined as:
+$f_{l}^{ij}=[\mathbf{f},S_{i,j}],$ (16)
+where $\mathbf{f}$ denotes a vector of available attributes related to the activation of the decentralized safety filter, and the binary indicator $S_{i,j}$ explicitly signals whether the safety filter has been activated between the two agents, thereby preventing setpoint update.
+IV-C Policy parameterization
+To demonstrate our approach we train the agents with a simple independent Q-Learning algorithm [16] and employ the same GNN architecture proposed in [8] for action-value estimation.
+The architecture integrates Graph Attention Network (GAT) [17] and Long Short-Term Memory (LSTM) layers [18], enabling effective modeling of spatial and temporal dependencies essential for collaborative decision-making.
+The GAT layers allow agents to dynamically attend to and integrate neighbor information, significantly improving inter-agent coordination by adaptively sharing latent representations through message passing.
+The LSTM captures temporal dependencies and partial observability, allowing agents to leverage historical context in their decisions.
+The architecture is trained using a centralized training decentralized execution (CTDE) framework, where the agents optimize the same action-value function parameterization during training while acting independently based on local observations and the interaction with the other agents.
+V Experimental Setup
+In this section, we describe in detail the specific environment implementation and training configurations employed for our MARL agents.
+V-A Dynamic network bridging environment implementation
+Environment: The simulated environment is normalized to a $[0,1]\times[0,1]$ grid, though agents can move beyond these limits.
+Each entity has a communication radius of 0.20, which defines the area within which it can detect and communicate with others.
+Agent observations: At each time step, agents receive an observation consisting of their local one-hop graph structure.
+This includes node-level features for themselves and all directly connected neighbors, as well as edge-level features for each neighbor pair.
+An example is shown in Figure 3, where Agent A observes Agent B.
+The observation includes both their node features $f_{n}^{A},f_{n}^{B}$ and the corresponding edge features $f_{i}^{AB}$ from a successful connection.
+Node-level features include the agent’s ID, current coordinates, last action taken, and the positions of the targets A and B.
+These features provide basic local context for decision-making.
+Edge-level features reflect spatial and safety-related information, as introduced in Section IV-B.
+For each neighboring agent $j$, agent $i$ observes relative position offsets $\Delta p_{x}=p_{x,i}-p_{x,j}$, $\Delta p_{y}=p_{y,i}-p_{y,j}$, and their Euclidean distance $d_{i,j}=\sqrt{\Delta p_{x}^{2}+\Delta p_{y}^{2}}$.
+Including safety-informed edge features allows agents to adjust their actions based not only on position but also on the state of the safety filter.
+This supports more informed and coordinated decision-making during training and execution.
+Agent actions: At the beginning of every time-step, agents make decentralized decisions, generating target points $w$, utilizing their local observation to choose the direction of their next movement.
+This is encoded in a two-dimensional action space, with each dimension having three options: move forward, move backward, or hold along the corresponding $p_{x}$ or $p_{y}$ axis.
+Given the direction chosen by the agents, a new target point $x_{sp}$ is calculated for each agent using a fixed offset equal for both the $p_{x}$ and $p_{y}$ axes.
+Reward: Agents are rewarded using a reward function that motivates agents to form a connected network bridging the targets. Such reward signal combines three components:
+Base connectivity reward: The base reward, $R_{\text{base}}$, encourages larger connected components:
+$R_{\text{base}}(s)=\frac{|C_{\text{max}}(s)|}{|\mathcal{V}|},$ (17)
+where $|C_{\text{max}}(s)|$ represents the size of the largest connected component in state $s$, and $|\mathcal{V}|$ is the total number of entities.
+Centroid distance penalty: The centroid distance penalty, $P_{\text{cent}}$, is based on agents’ positions relative to the targets, penalizing them based on the Euclidean distance between the centroid of the agents’ positions and the central point between the targets.
+Target path bonus: A bonus, $B_{\text{path}}=100$, is awarded if a path exists between the two targets.
+The overall reward combines these three elements:
+$R(s,a)=\begin{cases}B_{\text{path}}(s)&\mbox{if }\exists\mathrm{path}(T_{1},T_{2});\par\\ R_{\text{base}}(s)-P_{\text{cent}}(s),&\mbox{otherwise}.\end{cases}$ (18)
+Simulated UAV dynamics: For our experiments, we use a six-degrees-of-freedom (6-DOF) unmanned aerial vehicle (UAV) as an agent, where its dynamics, control laws, and safety sets used in our simulations are defined in the Appendix of [7].
+For the UAV simulations, the nonlinear model of the drone is used.
+Target mobility: To increase task complexity, each target moves deterministically toward waypoints located outside the communication range of any agent.
+When a target reaches a waypoint, it pauses for a fixed number of time-steps (5 in our implementation) before selecting a new one.
+Waypoint selection ensures that at least two agents (and up to $N$) are required to maintain a communication bridge between the targets.
+Figure 3: A typical configuration of our system.
+Agents are represented with letters A, B, C and targets with T.
+Agent A and B have established a successful connection.
+Agent A receives node features $f_{n}^{A}$, $f_{n}^{A}$ and edge features $f_{n}^{AB}$ as inputs.
+V-B Training configurations
+We train our agents using episodes of 100 time steps, for a total of one million steps.
+For generalization, we initialize agent positions, target locations, and target motion patterns using different seeded random states across training and testing scenarios.
+During evaluation, each trained agent is tested in 100 new episodes, and we measure performance metrics with the safety filter consistently enabled.
+We systematically evaluated different training configurations to assess the impact of our safety mechanism:
+  • Baseline: A vanilla MARL approach with no safety filter enabled. Edge features include only the base features $\mathbf{f}$.
+  • Approach A (Safety Filter): Our main contribution.
+Agents use Algorithm 1 to prevent safety infringements.
+The binary safety activation indicator is incorporated into the agents’ edge-level features ($\mathbf{f}$ and $S_{i,j}$).
+  • Approach B (Safety + Penalty): We add an explicit penalty of $-10$ to Approach A whenever an agent attempts an action that would trigger the safety filter.
+  • Approach C (Safety + Penalty + Early Truncation): Building on Approach B, we truncate episodes immediately upon safety-triggering situations, thus enforcing even stricter safety adherence.
+Every approach adopts the same learning algorithm and GNN architecture.
+In any configuration employing the safety filter, an agent violating (7) has its trajectory toward $x_{sp}$ blocked at its current position.
+Furthermore, we train an alternative of Approach A, called Approach A-Node, to investigate the benefits of our proposed integration of safety-informed edges IV-B.
+In this alternative approach, the agents utilize only the node-level features with the addition of a binary indicator $\mathbb{S}$ in case the agent had his safety filter activated during the previous step.
+VI Results
+(a) Mean return achieved during training. Approach A (with safety-informed edges) converges to higher rewards compared to Approach A-Node.
+(b) Average number of safety filter activations per episode. Approach A consistently triggers fewer interventions.
+Figure 4: Impact of safety-informed edge features.
+Table I summarizes our experimental evaluation of various learned MARL strategies under different training conditions.
+This comparison was conducted to evaluate the impact of our safety filter.
+We report the average communication coverage achieved by the agents and the average number of times the safety filter was activated during evaluation.
+Furthermore, Figure 4 focuses on a comparison between Approach A and Approach A-Node in terms of the mean return, i.e.
+the sum of the rewards, achieved by the agents during training and the mean number of safety activations triggered by the agents.
+VI-A Impact of the safety filter
+Baseline agents, trained without any safety mechanisms, achieved an average coverage of $(37.63\pm 16.62)\%$ but required frequent interventions from the safety filter, averaging $26.99\pm 12.31$ triggers per episode.
+In contrast, our Approach A, which incorporates our safety filter during training without explicit penalties, significantly improved average coverage to $(50.14\pm 12.72)\%$ and reduced safety filter activations to $20.45\pm 7.49$.
+Introducing explicit penalties alongside the safety mechanism (Approach B) maintained competitive coverage at $(45.02\pm 11.81)\%$ and further reduced the frequency of safety filter interventions to $15.25\pm 7.82$.
+However, implementing truncation upon safety violations (Approach C) considerably decreased task performance to $(10.86\pm 6.44)\%$, though it also minimized the safety triggers to $9.76\pm 5.72$.
+These results demonstrate that integrating our safety filter during training and including safety-informed edges allows agents to learn better strategies for the dynamic network bridging environment.
+Results on Approach B and Approach C highlight how the introduction of explicit penalties might reduce safety interventions, which could have practical benefits in specific applications.
+However, this led to a decrease in task performance, which was conspicuous for Approach C, suggesting that agents might exploit the activation of the safety filter to achieve better performance.
+VI-B Impact of the safety-informed edges
+Figure 5: Sequence of actions executed by agents trained with Approach A when deployed in a LVC environment.
+Figure 4(a) shows the mean return (cumulative reward) curves for Approach A and Approach A-Node, over the course of training.
+Approach A (with safety edge features) converges to higher returns compared to Approach A-Node, indicating that safety-informed edges facilitate more effective coordination strategies.
+Additionally, Figure 4(b) presents the average frequency of safety filter activations across training episodes.
+The results show that Approach A triggers fewer safety interventions than Approach A-Node, particularly in the later stages of training.
+These findings suggest that edge-level safety features, rather than node-level, help achieve better performance while incurring in less safety activations.
+Figure 5 shows an example of a sequence of actions operated by our decentralized agents trained with Approach A when deployed in a live, virtual, and constructive (LVC) environment.
+VII Discussion
+Our findings demonstrate that integrating a decentralized safety filter with MARL provides a versatile solution for multi-agent networked tasks, where agents can directly communicate with their one-hop neighbors.
+By embedding the safety mechanism in each agent, we ensure local checks on potential collisions, thereby containing safety concerns within these immediate communication ranges.
+This design preserves the decentralized nature of our approach while providing global safety assurances.
+We applied the framework to the dynamic network bridging task, which we extended to include more realistic target mobility and non-linear UAV dynamics.
+To support coordination, we included safety-informed edge features in the agents’ observations.
+These features indicate when the safety filter is active, allowing agents to adjust their actions based on local constraints.
+This helped agents avoid collisions, maintain communication links, and resolve potential deadlock situations through coordination.
+Empirical evaluations confirm that our approach effectively improves task performance while assuring safety.
+Although adding explicit penalties or early-termination strategies reduced safety activations, these methods often compromised the agent’s utility.
+In contrast, our decentralized filter and safety-informed edges guide safe exploration without requiring extensive reward engineering or outcome truncations.
+It is important to highlight that the design principles generalize naturally to larger numbers of agents and targets and it can be applied to different environments than the one under study here.
+For example, we can apply similar safety filters to other networked navigation tasks or expand the existing safety filter to include complex metrics.
+Looking ahead, we plan to investigate scalability for swarms of varying sizes, refine message-passing protocols for more sophisticated risk management, and evaluate domains with adversarial behaviors or dynamic constraints.
+Overall, our results highlight the potential of combining local safety controllers with MARL in networked multi-agent settings, underscoring that edge-level safety signals are crucial for guiding agents toward globally safe and high-performing outcomes.